# Periodic Chains Scheduling on Dedicated Resources - A Crucial Problem in Time-Sensitive Networks

### v1

@@ Abstract
Periodic messages transfer data from sensors to actuators in cars, planes, and complex production machines. When considering a given routing, the unicast message starts at its source and goes over several dedicated resources to reach its destination. Such unicast message can be represented as a chain of point-to-point communications. Thus, the scheduling of the periodic chains is a principal problem in time-triggered Ethernet, like IEEE 802.1Qbv Time-Sensitive Networks. This paper studies a strongly NP-hard periodic scheduling problem with harmonic periods, task chains, and dedicated resources. We analyze the problem on several levels of generality and complexity and provide the corresponding proofs. We describe a solution methodology to find a feasible schedule that minimizes the chains’ degeneracy related to start-to-end latency normalized in the number of periods. We use the local search with the first fit scheduling heuristic, which we warm-start with a constraint programming model. This notably improves the schedulability of instances with up to 100% utilization and thousands (and more) of tasks, with high-quality solutions found in minutes. An efficient constraint programming matheuristic significantly reduces the degeneracy of the found schedules even further. The method is evaluated on sets of industrial-, avionic-, and automotive-inspired instances.
@@
Proof of Strong NP-hardness
Constraint Programming Warm Start
Constraint Programming Matheuristic
@@ 1.1 Motivation
Ethernet is widely used in everyday life due to its high throughput and availability of SW stacks. Still, it lacks time determinism, essential for real-time applications, such as in-vehicle networks, industrial production lines, and audio streaming, that are very sensitive to delays. Time-sensitive networking (TSN) is a set of new standards that enables sending messages in a time-triggered manner. To use TSN efficiently and satisfy various requirements, many researchers [2, 3] have developed algorithms that can synthesize a schedule of periodic messages. Their results are versatile, but the network utilization and algorithm performance are often very low. Thus, in this paper, we omit a complex behavior of the TSN (queues, deadlines, frame isolation), and we concentrate on a crucial part of this problem: the scheduling of periodic tasks. We focus on harmonic periods of the messages (i.e., for each pair of distinct periods, it holds that the longer period is divisible by the shorter one) that are widely used in engineering practice and help with the tractability of the Periodic Scheduling Problem (PSP).
@@ 1.2 Related Work
Scheduling of both harmonic and non-harmonic periodic problems, as well as some special cases, was studied in [7, 8]. The authors presented a local schedulability condition for the general case of PSP. Motivated by the desire to increase the schedulability for highly-utilized instances, alternative views of the PSP based on bin-trees [9] and bin packing [10, 11, 12] were presented. These perspectives enabled the relevant authors to devise heuristics and exact methods to efficiently find feasible schedules.
The PSPs are often found in the domains of avionics and automotive. Building on their previous work, authors of [13] used Integer Linear Programming (ILP) to partition tasks on different processors of modern aircraft. In [14], the goal was to schedule tasks in the avionics domain to maximize a leeway between them. This way, the occasional increase in the processing time of the task would not interfere with the rest of the schedule. A game-theoretic approach to tackle a principally similar problem was presented in [15]. In [16], the authors tackled the scheduling of so-called partitions on shared processors. Their idea utilized a greedy heuristic, streamlining an otherwise exhaustive approach without a significant loss in the quality of the schedules. In [17], communication between modules of the aircraft is considered, and the authors solve it with constraint-generating ILP procedure.
In automotive domains, many papers tackle scheduling problems related to specific protocols and architectures. In [18], the authors schedule tasks for AUTOSAR architecture to eliminate load peaks in time. ILP and heuristics for the same architecture were also investigated in [19]. FlexRay, a protocol often coupled with the AUTOSAR specification, was studied in [10] and more recently in [20, 21]. In the latter, the authors consider multiple message variants so the same schedule can be repurposed for different vehicle models.
In the previous paragraphs, the authors mostly focused on partitioning tasks on different resources and finding appropriate schedules. However, in cases with messages passing from one resource to another, the assignment of the tasks to the resources is determined beforehand during the routing phase.
@@
Early work on offline scheduling for TTEthernet (an example of time-triggered ethernet) was presented in [24]. The authors studied several network topologies, including line topology, and used satisfiability modulo theories (SMT) solver in their proposed incremental workflow. Many papers emerged with the introduction of the TSN; see survey papers [2, 3]. In the latter, the authors categorized several different methods and performed experimental comparisons.
In [25], the authors developed an SMT model to shorten gate control lists controlling the schedules; their length is a limited resource in practice. Characteristics of the real industrial networks were exploited in [26] to develop a two-stage scheduling method; authors were able to isolate parts of the problem and solve them independently of each other. The TSN formalism can be further extended to include, e.g., authentication, as is demonstrated in [27]. Cryptography keys are disclosed with a time delay, enabled by the TSN, so nodes of the network may authenticate themselves. In our previous TSN papers [28, 29], concentrating on a complete set of TSN requirements and tight time windows, we achieved very low utilization of the communication resources.
On the other hand, in [30], where we presented a first-fit heuristic for a very similar problem (dedicated resources with soft constraints in the form of criteria instead of hard ones), we achieved nearly 100% utilization of the resource. This observation triggered the research presented in this paper, concentrating on the principal aspects of scheduling periodic chains on dedicated resources.
Apart from communications, PSPs apply to manufacturing and maintenance. In [31, 32], the scheduling problems with (almost) periodic maintenance windows are studied; there, the maintenance task is frequently repeated, while the manufacturing tasks are essentially aperiodic. This framework is useful in practice since inspecting the machines periodically reduces the risk of failures. A practical application can be found in [33], where the authors schedule patients’ non-preemptive treatments around the medical machine’s maintenance.
@@ 1.3 Outline and Contributions
In this work, we focus on the scheduling of messages, represented by chains of non-preemptive tasks. Each task is to be performed on a dedicated resource, and all tasks of the chain have the same period from a set of harmonic periods. We primarily optimize the degeneracy, which is related to the message’s start-to-end latency normalized in the number of periods. We build upon our single resource scheduling paper [12] and extend our heuristic outlined in [30].
In Section 2, we formally describe the problem and its structural properties, including the connection between PSP with harmonic periods and bin-packing. In Section 3, we discuss the complexity of several special cases of PSPs. We prove it is polynomial for a single period, whereas it is strongly NP-Hard for two periods. In Section 4, we describe a heuristic algorithm and Constraint Programming (CP) approaches to warm-start the algorithm and fine-tune the final solution. In Section 5, we discuss the experimental results, and in Section 6, the conclusions are drawn.
@@
  • 3.
    CP-based matheuristic improving found schedule
2 Problem Description
2.1 Problem Statement
Tasks and Resources
@@ 2.1 Problem Statement
We consider a set of $\kappa$ chains. A chain $\mathcal{C}^{k}$ with $k\in\{1,\ldots,\kappa\}$ is composed of tasks $J_{i}^{k}$ $i\in\{1,\ldots,n^{k}\}$ with integer processing times $p_{i}^{k}$. Each task is to be performed on a pre-defined unit capacity resource among $\{1,\ldots,M\}$. The index of the resource processing $J_{i}^{k}$ is denoted by $\pi_{i}^{k}$.
In TSN applications, a chain corresponds to a message sent periodically, and resources are successive communication links used to deliver the message. A point-to-point communication performed by such communication link corresponds to a task. We elaborate on this correspondence in Fig. 1(a). Notice that since the resource corresponds to a communication link rather than a node, the messages in the example consist of three point-to-point communications (i.e., tasks) but travel between four nodes of the network. From a scheduling point of view, a resource can be viewed as a machine. Our problem thus shares some features with job-shop problems.
@@
The tasks of a chain are periodic and share the same period. We assume that the periods belong to a harmonic set $\tau(1),\ldots,\tau(\rho)$ of size $\rho$, so that $\forall l\,\in\nolinebreak\{2,\ldots,\rho\},\tau(l)=b_{l}\tau(l-1)$ where $b_{l}$ is an integer. We define the least period $w=\tau(1)$, and we define the hyperperiod $\hat{H}$ to be the least common multiple of periods. Due to the harmonic nature of the period set, $\hat{H}$ is equal to the longest period $\tau(\rho)$. All of the tasks of chain $\mathcal{C}^{k}$ have period $T^{k}=\tau(\beta^{k})$, where $\beta^{k}$ is the index of period $T^{k}$ in the ordered set of periods.
For resource $m$ we denote by $\mathcal{J}_{m}$ the set of tasks $J_{i}^{k}$ assigned to it (i.e. such that $\pi_{i}^{k}=m$). We define the utilization of resource $m\in\{1,\ldots,M\}$ as:
$U_{m}=\sum_{J_{i}^{k}\in\mathcal{J}_{m}}\frac{p_{i}^{k}}{T^{k}}$
The utilization $U_{m}\leq 1$ represents the percentage of time resource $m$ is busy in any feasible schedule. The utilization of an instance is the maximum utilization of its resources.
Notation
According to the standard Graham notation extended to handle PSPs in [4], our problem is denoted by ${PD|T_{i}^{harm},jit=0,chains|criteria}$. $PD$ stands for parallel dedicated processors, $chains$ for chains, $T_{i}^{harm}$ for harmonic periods, $jit=0$ since no jitter is allowed (i.e., the starting time of a task is entirely determined by the starting time of its first occurrence and by its period). We assume that all tasks of a chain have the same period.
(a) Ten messages are sent from the leftmost node A to the rightmost node D via three communication links (i.e., resources). Corresponding chains $\mathcal{C}^{1}-\mathcal{C}^{10}$ consisting of three tasks $J_{i}^{k}$ are to be scheduled on resources 1, 2, 3 (assignment of tasks to resources $\pi_{i}^{k}=i$ is shown by dotted ellipses). For simplicity, no messages are sent via links in the opposite direction 4, 5, 6.
(b) Schedule of 10 chains on 3 resources spanning over one hyperperiod. Tasks of the chain have the same color. Total degeneracy is 2. Start-to-end latency $S^{k}$ is shown for cyan $\mathcal{C}^{1}$ and yellow $\mathcal{C}^{3}$ chains, which both have degeneracy $D^{k}=1$.
Figure 1: Illustrative communication case and associated schedules for instance with periods $\tau(1)=40,\tau(2)=80,\tau(3)=240$. 10 chains are to be scheduled; $\mathcal{C}^{1}-\mathcal{C}^{2}$ have period 40 (thus $T^{1}=T^{2}=\tau(1)$, meaning $\beta^{1}=\beta^{2}=1$), $\mathcal{C}^{3}-\mathcal{C}^{7}$ have period 80, and $\mathcal{C}^{8}-\mathcal{C}^{10}$ have period 240. Notice that all the chains are sent from left to right in this example; communication links in the other direction are not used. Utilization $U_{m}=0.96$ holds for all resources.
Schedules
@@
A schedule assigns an integer start time $t_{i}^{k}$ to the first occurrence of each task $J_{i}^{k}$, such that the precedences are fulfilled:
$t_{i+1}^{k}\geq t_{i}^{k}+p_{i}^{k}\quad\forall k\in\{1,\ldots,\kappa\},\forall i\in\{1,\ldots,n^{k}-1\},\quad$ (1)
and at any time instant $s$, a processor performs at most one task:
@@
Notice that if independent periodic tasks $\mathcal{J}_{m}$ can be scheduled on their respective resources, we can derive a schedule that satisfies precedence constraints as well. Indeed, assume we are given a feasible schedule $\sigma$ of the independent periodic tasks where $\sigma_{i}^{k}<T^{k}$ is the start time of the first occurrence of $J_{i}^{k}\in\nolinebreak\mathcal{J}_{m}$. Then, we can satisfy the precedence constraints (1) of each chain $k$ by postponing a subsequent task in the chain to the next period whenever needed to satisfy the precedence constraint. Formally, we can set:
$t_{i}^{k}=\sigma_{i}^{k}+q_{i}T^{k}$
with $q_{1}=0$. If $\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}$, then we set $q_{i+1}=q_{i}$. If $\sigma_{i+1}^{k}<\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}+T^{k}$ we set $q_{i+1}=q_{i}+1$. Otherwise we set $q_{i+1}=q_{i}+2$. Observe that in the latter case, we know that $\sigma_{i}^{k}+p_{i}^{k}<2\cdot T^{k}\leq\sigma_{i+1}^{k}+2\cdot T^{k}$. So, in any case, $t_{i}^{k}+p_{i}^{k}\leq t_{i+1}^{k}$ so that the precedence constraints are satisfied.
Conversely, a feasible integer schedule of the chains induces a feasible schedule $\sigma$ of independent periodic tasks of $\mathcal{J}_{m}$ on resource ${m}$ by setting
$\sigma_{i}^{k}=t_{i}^{k}\mod T^{k}.$
with the additional property that $\sigma_{i}^{k}<T^{k}$. This schedule is called the core schedule of resource $m$.
Such schedules, shown within a single hyperperiod, can be illustrated as in Fig. 1. There, 10 chains are delivered using a sequence of resources 1, 2, and 3. Precedence constraints for chains hold, and no overlaps occur on either resource.
Criteria
All criteria we consider are related to the start-to-end latency of chains, i.e., the delay between the moment when the chain’s first task started and the chain’s last task ended. In communications, this refers to the delay between the moment the message was sent from the first (source) resource and when the message was processed at the last (sink) resource.
Let $t$ be a schedule. We define the following criteria:
Start-to-end latency of chain $k$:
$S^{k}=t_{n^{k}}^{k}+p_{n^{k}}^{k}-t_{1}^{k}$
Degeneracy of chain $k$:

### v2

@@ Abstract
Periodic messages transfer data from sensors to actuators in cars, planes, and complex production machines. When considering a given routing, the unicast message starts at its source and goes over several dedicated resources to reach its destination. Such unicast message can be represented as a chain of point-to-point communications. Thus, the scheduling of the periodic chains is a principal problem in time-triggered Ethernet, like IEEE 802.1Qbv Time-Sensitive Networks. This paper studies a strongly NP-hard periodic scheduling problem with harmonic periods, task chains, and dedicated resources. We analyze the problem on several levels and provide proofs of complexity and approximation algorithms for several special cases. We describe a solution methodology to find a feasible schedule that minimizes the chains’ degeneracy related to start-to-end latency normalized in the number of periods. We use the local search with the first fit scheduling heuristic, which we warm-start with a constraint programming model. This notably improves the schedulability of instances with up to 100% utilization and thousands (and more) of tasks, with high-quality solutions found in minutes. An efficient constraint programming matheuristic significantly reduces the degeneracy of the found schedules even further. The method is evaluated on sets of industrial-, avionic-, and automotive-inspired instances.
@@
Offset schedules guarantee worst-case degeneracy for acyclic instances
Constraint programming successfully initializes the local search
Per-period matheuristic efficiently improves existing schedules
Publicly available diverse set of PSP instances
@@ 1.1 Motivation
Ethernet is widely used in everyday life due to its high throughput and availability of SW stacks. Still, it lacks time determinism, essential for real-time applications, such as in-vehicle networks, industrial production lines, and audio streaming, that are very sensitive to delays. Time-sensitive networking (TSN) is a set of new standards that enables sending messages in a time-triggered manner. To use TSN efficiently and satisfy various requirements, many researchers [2, 3] have developed algorithms that can synthesize a schedule of periodic messages. Their results are versatile, but the network utilization and algorithm performance are often very low. Thus, in this paper, we omit a complex behavior of the TSN (link delays, queues, deadlines, frame isolation, see [4]), and we concentrate on a crucial part of this problem: the scheduling of periodic tasks. We focus on harmonic periods of the messages (i.e., for each pair of distinct periods, it holds that the longer period is divisible by the shorter one) that are widely used in engineering practice and help with the tractability of Periodic Scheduling Problem (PSP).
@@ 1.2 Related Work
Scheduling of both harmonic and non-harmonic periodic problems and several special cases were studied in [8, 9]. The authors presented a local schedulability condition for the general case of PSP. Motivated by the desire to increase the schedulability for highly-utilized instances, alternative views of PSPs based on bin-trees [10] and bin packing [11, 12, 13] were presented. These perspectives enabled the relevant authors to devise heuristics and exact methods to find feasible schedules efficiently.
PSPs are often found in the domains of avionics and automotive. Building on their previous work, authors of [14] used Integer Linear Programming (ILP) to partition tasks on different processors of modern aircraft. In [15], the goal was to schedule tasks in the avionics domain to maximize a leeway between them. This way, the occasional increase in the processing time of the task would not interfere with the rest of the schedule. A game-theoretic approach to tackling a similar problem was presented in [16]. In [17], the authors tackled the scheduling of so-called partitions on shared processors. Their idea utilized a greedy heuristic, streamlining an otherwise exhaustive approach without a significant loss in the quality of the schedules. In [18], communication between modules of the aircraft is considered, and the authors solve it with constraint-generating ILP procedure.
In the automotive domain, many papers tackle scheduling problems related to specific protocols and architectures. In [19], the authors schedule tasks for AUTOSAR architecture to eliminate load peaks. ILP and heuristics for the same architecture were also investigated in [20]. FlexRay, a protocol often coupled with the AUTOSAR specification, was studied in [11] and more recently in [21, 22]. In the latter, the authors consider multiple message variants so the same schedule can be repurposed for different vehicle models.
Apart from communications, PSPs apply to manufacturing and maintenance. In [23, 24], the scheduling problems with (almost) periodic maintenance windows are studied; there, the maintenance task is frequently repeated, while the manufacturing tasks are essentially aperiodic. This framework is helpful in practice since regularly inspecting the machines reduces the risk of failures. A practical application can be found in [25], where the authors schedule patients’ non-preemptive treatments around the medical machine’s maintenance.
The rest of this section focuses on papers regarding PSP with dedicated resources, which is the main focus of this paper. Messages are transmitted through the network, and the tasks are assigned to resources beforehand, given the network’s routing.
@@
Another domain is scheduling for cloud radio access networks. In [28, 29], the authors focus on periodic scheduling with a single period and uniform processing times of the messages. Especially in the latter, the no-buffering constraint complicates the problem. The authors provide a polynomial scheduling algorithm applicable, when the network load is sufficiently low, and a general greedy method used in other cases.
Early work on offline scheduling for TTEthernet (an example of time-triggered ethernet) was presented in [30]. The authors studied several network topologies, including line topology, and used satisfiability modulo theories (SMT) solver in their proposed incremental workflow. Many papers emerged with the introduction of the TSN; see survey papers [2, 3]. The authors categorized several methods and performed experimental comparisons in the latter.
In [31], the authors developed an SMT model to shorten gate control lists controlling the schedules; their length is a limited resource in practice. Characteristics of the real industrial networks were exploited in [32] to develop a two-stage scheduling method; authors were able to isolate parts of the problem and solve them independently. The TSN formalism can be further extended to include, e.g., authentication, as is demonstrated in [33]. The cryptography keys are disclosed with a time delay (enabled by the TSN) so that the network nodes can authenticate themselves.
In [34, 4], the authors focused on a complete set of TSN requirements and tight time windows and achieved only low utilization of communication resources. On the other hand, in [35], a first-fit heuristic for a very similar problem was presented (dedicated resources with soft constraints instead of hard ones), and the authors achieved nearly 100% utilization of the resource. This observation triggered the research presented in this paper, concentrating on the principal aspects of scheduling periodic chains on dedicated resources.
@@ 1.3 Outline and Contributions
In this work, we primarily focus on the scheduling of messages, represented by chains of non-preemptive tasks. Each task is to be performed on a dedicated resource, and all tasks of the chain have the same period from a set of harmonic periods. We optimize the degeneracy, which is related to the message’s start-to-end latency normalized in the number of periods. We build on a single-resource scheduling paper [13] and extend the heuristic outlined in [35].
In Section 2, we formally describe the problem and its structural properties, including the connection between PSP with harmonic periods and bin-packing. In Section 3, we discuss the complexity of several special cases of PSPs. In Section 4, we describe a heuristic algorithm and Constraint Programming (CP) approaches to warm-start the algorithm and fine-tune the final solution. In Section 5, we discuss the experimental results, and in Section 6, the conclusions are drawn. The list of the most used symbols follows. Furthermore, instances used in our experiments are available [36].
@@
    Guarantees of maximum degeneracy for instances with acyclic support graphs and a bottleneck resource
  • 3.
@@
  • 4.
    CP-based matheuristic for improving existing schedule
2 Periodic Scheduling with Dedicated Resources
2.1 Model Abstraction
In this paper, we use concepts of tasks $J$, chains of tasks $\mathcal{C}$ and resources (described in detail in Section 2.2) which can be used to model diverse real-life systems and problems. In periodic manufacturing, the resources would correspond to production machines, chains to orders that need to be processed by several machines, and tasks to operations of the order.
Specific hardware and software of the communications network influence how the model of the associated PSP is created. For example, when the CPU of the network node performs time-consuming processing on each message, then the CPU itself is the resource since it is the communication bottleneck. For example, in [37, 29], tasks correspond to computations in the network nodes, and passing a message through a link induces a precedence delay between the tasks.
In modern TSN applications, the network nodes use massive parallel hardware. Thus, the network node itself is not the communication bottleneck. Therefore, the resources correspond to communications links between network nodes. Chains correspond to messages sent from one network node to another, and the task is a point-to-point communication via a single link. We elaborate on the relationship between the network topology and associated PSP in further detail in Fig. 2(a) with four network nodes A, B, C, and D depicted by blue rectangles, and resources 1-6 shown as blue arrows with ellipses.
Our model and benchmarks also capture other realistic phenomena. For example, in automotive, zonal architecture is often used as a communication architecture, where the peripheral links connected to sensors have much slower speeds than the links connecting central nodes. This is modeled by proportionally increasing the processing times of tasks dedicated to slower links.
2.2 Problem Description
2.2.1 Tasks, Resources
Consider a set of $\mu$ chains of tasks. Each chain corresponds to a message sent periodically. In the rest of the paper superscript $k$ is used to index the chain, while subscripts are used to describe the task number within the chain. So, a superscript does not indicate a power function.
Definition 1 (Tasks, chains, and resources).
@@ 2.2.1 Tasks, Resources
A chain $\mathcal{C}^{k}$ with $k\in\{1,\ldots,\mu\}$ is composed of tasks $J_{i}^{k}$, $i\in\{1,\ldots,n^{k}\}$ with integer processing times $p_{i}^{k}$. The tasks of a chain $\mathcal{C}^{k}$ are periodic and share the same period $T^{k}$. Once a task $J_{i}^{k}$ is started, it is repeated every $T^{k}$ time unit. Each task and its repetitions are to be performed on a pre-defined unit capacity resource $\{1,\ldots,M\}$. The index of the resource processing $J_{i}^{k}$ is denoted by $\pi_{i}^{k}$. For a resource $m\in\{1,\ldots,M\}$ we denote by $\mathcal{J}_{m}$ the set of tasks $J_{i}^{k}$ assigned to it (i.e. such that $\pi_{i}^{k}=m$).
The structure of an instance will be analyzed through the notion of a support graph:
Definition 2 (Support graph).
We call the support graph of an instance the directed graph $G=(V,E)$ where $V=\{1,\ldots,M\}$ represents the resources, and there is an arc from node $m$ to node $m^{\prime}$ if there is a chain $\mathcal{C}^{k}$ such that two consecutive tasks of the chain $J_{i}^{k},J_{i+1}^{k}$ satisfy $\pi_{i}^{k}=m,\pi_{i+1}^{k}=m^{\prime}$.
Examples of such support graphs are shown in Fig. 1, including the chains which induced the graphs’ edges.
(a) Support graph with three vertices and without directed cycles.
(b) Support graph with four vertices and with directed cycles produced by chains $\mathcal{C}^{2},\mathcal{C}^{3}$.
Figure 1: Examples of support graphs. Their vertices (i.e., resources) are the dashed ellipses, and their edges are shown as dashed arrows. The green chains and green tasks induced the edges of the support graphs.
2.2.2 Harmonic Periods, Feasible Schedules
In this paper, we assume that all periods belong to a harmonic set:
Definition 3 (Harmonic set of periods).
For any chain $\mathcal{C}^{k}$, its period $T^{k}$ is in an harmonic set of integers $\{\tau(1),\ldots,\tau(\rho)\}$. Such a set satisfies
$\forall l\,\in\nolinebreak\{2,\ldots,\rho\},\tau(l)=b_{l}\tau(l-1)$ (1)
where $b_{l}$ is an integer. The least period is also denoted $w=\tau(1)$, and we define the hyperperiod
$\hat{H}=lcm(\tau(1),\ldots,\tau(\rho))=\tau(\rho)$ (2)
@@ 2.2.2 Harmonic Periods, Feasible Schedules
We assume that all of the tasks of chain $\mathcal{C}^{k}$ have the same period $T^{k}=\tau(\beta^{k})$, where $\beta^{k}$ is the index of period $T^{k}$ in the ordered set of periods.
Definition 4 (Schedule).
@@
A schedule assigns an integer start time $t_{i}^{k}$ to the first occurrence of each task $J_{i}^{k}$. Moreover, an occurrence of $J_{i}^{k}$ is performed at each time $t_{i}^{k}+\alpha T^{k}$, for any integer $\alpha\geq 0$.
A schedule is feasible if it satisfies precedence constraints and resource constraints defined below:
Definition 5 (Precedence constraints).
Task $J_{i}^{k}$ precedes task $J_{i+1}^{k}$:
$t_{i+1}^{k}\geq t_{i}^{k}+p_{i}^{k}\quad\forall k\in\{1,\ldots,\mu\},\forall i\in\{1,\ldots,n^{k}-1\},\quad$ (3)
Definition 6 (Resource constraints).
At any time instant $s$, a resource $m$ performs at most one task:
@@
Definition 7 (Utilization).
We define the utilization of resource $m\in\{1,\ldots,M\}$ as:
$U_{m}=\sum_{J_{i}^{k}\in\mathcal{J}_{m}}\frac{p_{i}^{k}}{T^{k}}$ (5)
The utilization of an instance is the maximum utilization of its resources.
A necessary condition of the feasibility of a set of periodic tasks on a resource $m$ is that the utilization, i.e., the percentage of time resource $m$ is busy, should not exceed $1$. The fact that utilization is not greater than 1 implies that for each task $p_{i}^{k}<T^{k}$ (otherwise the task $J_{i}^{k}$ would be performed without any other task on its machine).
2.2.3 Objectives
All objective functions we consider are related to the start-to-end latency of chains, that is, the delay between the moment when the chain’s first task started and the chain’s last task ended. In communications, this refers to the delay between the moment the message was sent from the first (source) network node and is being processed by the first link (resource) and when the message was delivered via the last link to the last (sink) network node. However, minimizing the start-to-end latency of chains might lead to unnecessarily difficult problems. Previous work [38] has shown that measuring the number of periods crossed by a message is significant enough in practice and simplifies the problem’s complexity. So, we consider period-normalized objectives as in [35]. Moreover, discrete-time control of cyber-physical systems naturally matches such objectives. The controller operates with its sampling period [39], and if the message arrives even slightly after the end of the sampling period, it will need to wait until the end of the next sampling period. There is no gain in arriving earlier within the sampling period.
Let $t$ be a schedule. We define the following functions:
Definition 8 (Start-to-end latency of chain $k$).
$S^{k}\coloneqq t_{n^{k}}^{k}+p_{n^{k}}^{k}-t_{1}^{k}$ (6)
Definition 9 (Degeneracy of chain $k$).
$D^{k}\coloneqq\left\lceil\frac{S^{k}}{T^{k}}\right\rceil-1.$ (7)